For 0<x<5 and t>0 , solve the one dimensional heat flow equation Thita u / Thita t = 4 thita^2 u / Thita x^2 satisfying the condition u(t,0) = u(t,5) = 0 , u(0,x) = x
1
Expert's answer
2020-03-11T14:21:06-0400

We use the method of separation of variables


"u(x,t)=X(x)T(t)\to\left\{\begin{array}{l}
\displaystyle\frac{\partial u}{\partial t}=\displaystyle\frac{\partial\left(X(x)T(t)\right)}{\partial t}=X(x)T'(t)\\[0.3cm]
\displaystyle\frac{\partial^2 u}{\partial x^2}=\displaystyle\frac{\partial^2\left(X(x)T(t)\right)}{\partial x^2}=X''(x)T(t)
\end{array}\right."

Then,


"\frac{\partial u}{\partial t}=4\cdot\frac{\partial^2 u}{\partial x^2}\to\left.X(x)T'(t)=4X''(x)T(t)\right|\div\left(4X(x)T(t)\right)\\[0.3cm]
\frac{T'(t)}{4T(t)}=\frac{X''(x)}{X(x)}=-\lambda\to
\left\{\begin{array}{l}
X''(x)+\lambda X(x)=0\\[0.3cm]
T'(t)+4\lambda T(t)=0
\end{array}\right."

The boundary conditions take the form

"\left\{\begin{array}{l}
u(0,t)=X(0)T(t)=0\to\boxed{X(0)=0}\\[0.3cm]
u(5,t)=X(5)T(t)=0\to\boxed{X(5)=0}
\end{array}\right."



First, we solve the Sturm-Liouville problem for the function "X(x)"

In our case,

"\left\{\begin{array}{l}
X''(x)+\lambda X(x)=0\\[0.1cm]
X(0)=0\\[0.1cm]
X(5)=0
\end{array}\right."

We are looking for a solution in the form


"X(x)=e^{kx}\to X''(x)=k^2\cdot e^{kx}\to\\[0.3cm]
k^2\cdot e^{kx}+\lambda\cdot e^{kx}=0\to e^{kx}\cdot\left(k^2+\lambda\right)=0\to k^2=-\lambda\\[0.3cm]
\left[\begin{array}{l}
k_1=\sqrt{-\lambda}=i\sqrt{\lambda}\\[0.1cm]
k_2=-\sqrt{-\lambda}=-i\sqrt{\lambda}
\end{array}\right."

Conclusion,


"X(x)=A_1e^{i\sqrt{\lambda}x}+A_2e^{-i\sqrt{\lambda}x}\equiv
C_1\cos\left(\sqrt{\lambda}x\right)+C_2\sin\left(\sqrt{\lambda}x\right)\\[0.1cm]
X(0)=0=C_1\cos\left(\sqrt{\lambda}0\right)+C_2\sin\left(\sqrt{\lambda}0\right)\\[0.1cm]
\boxed{C_1=0}\\[0.1cm]
X(5)=0=C_2\sin\left(5\sqrt{\lambda}\right)\to5\sqrt{\lambda}=n\pi,n\in\mathbb{N}\\[0.1cm]
\boxed{\lambda_n=\left(\frac{n\pi}{5}\right)^2,n\in\mathbb{N}}"

Then,


"\boxed{X_n(x)=C_n\cdot\sin\left(\frac{n\pi x}{5}\right),n\in\mathbb{N}}"

For the function "T(t)" it is necessary to solve the equation


"T'(t)+\lambda_nT(t)=0\to\frac{dT}{dt}=-\lambda T(t)\to\frac{dT}{T}=-\lambda_n dt\\[0.1cm]
\ln|T(t)|=-\lambda_nt+\ln|A_n|\to\boxed{T_n(t)=A_n\cdot e^{-\lambda_n t}}"

Then, a particular solution to the equation has the form


"u_n(x,t)=X_n(x)T_n(t)=A_n\cdot e^{-\lambda_nt}\cdot C_n\cdot\sin\left(\frac{n\pi x}{5}\right)"

And the general solution of the equation has the form


"\boxed{u(x,t)=\sum\limits_{n=1}^{\infty}u_n(x,t)=\sum\limits_{n=1}^{\infty}B_ne^{-\lambda_nt}\sin\left(\frac{n\pi x}{5}\right)}"



It remains to find an expression for the coefficients "B_n". To do this, we use the initial condition, as well as the condition of orthogonality of the function of the Sturm-Liouville problem (without proof)


"\int\limits_0^5\sin\left(\frac{n\pi x}{5}\right)\cdot\sin\left(\frac{n\pi x}{5}\right)dx=\frac{5}{2}\\[0.3cm]
\int\limits_0^5\sin\left(\frac{n\pi x}{5}\right)\cdot\sin\left(\frac{m\pi x}{5}\right)dx=0\\[0.3cm]
\boxed{\int\limits_0^5\sin\left(\frac{n\pi x}{5}\right)\cdot\sin\left(\frac{m\pi x}{5}\right)dx=\frac{5}{2}\delta_{nm}}"

Then,


"u(x,0)=x=\sum\limits_{n=1}^{\infty}B_n\sin\left(\frac{n\pi x}{5}\right)\to\\[0.3cm]
\int\limits_0^5\times\left|\sum\limits_{n=1}^{\infty}B_n\sin\left(\frac{n\pi x}{5}\right)=x\right|\times\sin\left(\frac{m\pi x}{5}\right)dx\to\\[0.3cm]
\sum\limits_{n=1}^{\infty}B_n\int\limits_0^5\sin\left(\frac{n\pi x}{5}\right)\sin\left(\frac{m\pi x}{5}\right)dx=\int\limits_0^5 x\sin\left(\frac{m\pi x}{5}\right)dx\to\\[0.3cm]
\sum\limits_{n=1}^{\infty}B_n\delta_{nm}\frac{5}{2}=\int\limits_0^5 x\sin\left(\frac{m\pi x}{5}\right)dx\to\\[0.3cm]
B_m\frac{5}{2}=\int\limits_0^5 x\sin\left(\frac{m\pi x}{5}\right)dx"

It remains to calculate the indicated integral. To do this, we use the method of integration by parts


"\int\limits_0^5 x\sin\left(\frac{m\pi x}{5}\right)dx=\\[0.3cm]
=\left.\frac{-5x}{m\pi}\cdot\cos\left(\frac{m\pi x}{5}\right)\right|_0^5-\int\limits_0^5\left(-\frac{5}{m\pi}\cos\left(\frac{m\pi x}{5}\right)\right)dx=\\[0.3cm]
=-\frac{25}{m\pi}\cdot\cos(m\pi)+\left.\frac{25}{m^2\pi^2}\cdot\sin\left(\frac{m\pi x}{5}\right)\right|_0^5\to\\[0.3cm]
\boxed{\int\limits_0^5 x\sin\left(\frac{m\pi x}{5}\right)dx=-\frac{25\cdot(-1)^m}{m\pi}}"

Conclusion,


"B_m\frac{5}{2}=\frac{25\cdot(-1)^{m+1}}{m\pi}\to\boxed{B_m=\frac{10\cdot(-1)^{m+1}}{m\pi}}"

ANSWER

"u(x,t)=\sum\limits_{n=1}^{\infty}\left(\frac{10\cdot(-1)^{n+1}}{n\pi}\right)e^{-\lambda_nt}\sin\left(\frac{n\pi x}{5}\right)"

"\lambda_n=\left(\frac{n\pi}{5}\right)^2,n\in\mathbb{N}"